# Progressive Autoregressive Video Diffusion Models

### v1

@@ 4.2 Long video generation
Figure 4: Qualitative comparison of PA-M (ours), RW-M, PA-O-base (ours), RN-O-base, StreamingSVD from StreamingT2V (Henschel et al., 2024), and SVD-XT from Stable Video Diffusion (Blattmann et al., 2023a). Frames are evenly sampled from 1 minute long generated video, i.e. at 10, 20, 30, 40, 50, and 60 seconds. Our models can autoregressively generate 60-second, 1440-frame videos without quality degradation.
We also show strength of our method with qualitative comparison results in Fig. 4. Both of our models demonstrate strong performance in terms of frame fidelity and motion realism (e.g. camera motion, wave motion, and running gestures). PA-M outperforms PA-O due to additional fine-tuning with our proposed progressive noise levels Sec. 3.3, whereas PA-O simply inherits the pre-trained weights from Open-Sora v1.2 base model (Zheng et al., 2024). In contrast, SVD-XT (Blattmann et al., 2023a) shows severe artifacts that decreases frame validity, and StreamingSVD from StreamingT2V (Henschel et al., 2024) suffers from cumulative errors, resulting in degraded video quality as the sequence length increases. For more qualitative results, please refer to our website https://desaixie.github.io/pa-vdm/.
Full
Ablation 1
Ablation 2
Figure 5: Qualitative comparison for ablation study. Full represents for our full solution based on M-PA, Ablation 1 is with “chunk-by-chunk denoising” but without “keeping the clean frames available in temporal self-attention”. Ablation 2 is without both techniques. Frames are evenly sampled from 16-second generated videos.
4.3 Ablation Study
We conduct ablation studies on the PA-M model to evaluate the impact of “chunk-by-chunk denoising” and “keeping clean frames available in temporal self-attention”, as described in Sec. 4.1. Qualitative comparison has been shown in Fig. 5. In Ablation 1, we observe that the absence of clean frames in the input sequence prevents noisy frames from attending to previous clean frames, resulting in poor performance over a long duration. This also causes frame-to-frame discontinuity, which is more noticeable in the supplementary anonymous webpage. In Ablation 2, not decoding the video chunk-by-chunk leads to severe cumulative errors, causing the video to diverge after only a few seconds.
5 Related Works
The field of long video generation has faced significant challenges due to the computational complexity and resource constraints associated with training models on longer videos. As a result, most existing text-to-video diffusion models (Guo et al., 2023; Ho et al., 2022a; b; Blattmann et al., 2023a) have been limited to generating fixed-size video clips, which leads to noticeable degradation in quality when attempting to generate longer videos. Recent works are proposed to address these challenges through innovative approaches that either extend existing models or introduce novel architectures and fusion methods.
Freenoise (Qiu et al., 2024) utilizes sliding window temporal attention to ensure smooth transitions between video clips but falls short in maintaining global consistency across long video sequences. Gen-L-video (Wang et al., 2023), on the other hand, decomposes long videos into multiple short segments, decodes them in parallel using short video generation models, and later applies an optimization step to align the overlapping regions for continuity. FreeLong (Lu et al., 2024) introduces a sophisticated approach which balances the frequency distribution of long video features in different frequency during the denoising process. Vid-GPT (Gao et al., 2024) introduces GPT-style autoregressive causal generation for long videos.
More recently, Short-to-Long (S2L) approaches are proposed, where correlated short videos are firstly generated and then smoothly transit in-between to form coherent long videos. StreamingT2V (Henschel et al., 2024) adopts this strategy by introducing the conditional attention and appearance preservation modules to capture content information from previous frames, ensuring consistency with the starting frames. It further enhances the visual coherence by blending shared noisy frames in overlapping regions, similar to the approach used by SEINE (Chen et al., 2023). NUWA-XL (Yin et al., 2023) leverages a hierarchical diffusion model to generate long videos using a coarse-to-fine approach, progressing from sparse key frames to denser intermediate frames. However, it has only been evaluated on a cartoon video dataset rather than natural videos. VideoTetris (Tian et al., 2024b) introduces decomposing prompts temporally and leveraging a spatio-temporal composing module for compositional video generation.
Another line of research focuses on controllable video generation (Zhuang et al., 2024; Tian et al., 2024a; Hu, 2024; Zhu et al., 2024) and has proposed solutions for long video generation using overlapped window frames. These approaches condition diffusion models using both frames from previous windows and signals from the current window. While these methods demonstrate promising results in maintaining consistent appearances and motions, they are limited to their specific application domains which relies heavily on strong conditional inputs.
6 Discussion
In this work, we target long video generation, a fundamental challenge of current video diffusion models. We show that they can be naturally adapted to become progressive autoregressive video diffusion models without changing the architectures. With our progressive noise levels and the autoregressive video denoising process (Secs. 3.1 and 3.2), we obtain state-of-the-art results on long video generation at 1-minute long. Since our method does not require changing the model architectures, it can be seamlessly combined with many orthogonal works, paving the way for generating longer videos at higher quality, long-term dependency, and controllability.
A limitation of our method is the demand of a well-trained base video diffusion model. Similar to the replacement methods (Ho et al., 2022b; Zheng et al., 2024) and other approaches like StreamingT2V (Henschel et al., 2024), our method autoregressively applies a video diffusion model to generate long videos. Such autoregressive video generation poses huge challenge on the base video diffusion model. Some slight errors remaining in the “clean” latents ${\mathbf{x}}^{0}$ may not be noticeable in a single video clip; however, in the autoregressive scenario, these error can be carried onto later frames, resulting in quality degradation. Further more, as the video diffusion model is only trained on denoising latents of real video data, it may poorly handle such distribution shift towards the generated erroneous latents (Xie et al., 2024; Fan et al., 2024), resulting in more severe quality drop. This means that even after finetuning on our progressive noise levels, our method could still generate videos with some degree of quality degradation over time if the base video diffusion model is not well trained.

### v2

@@ 4.2 Long video generation
FIFO
Figure 5: Qualitative comparison of PA-M (ours), RW-M, PA-O-base (ours), RN-O-base, StreamingSVD from StreamingT2V [12], SVD-XT from Stable Video Diffusion [1], and FIFO-Diffusion [19]. Frames are evenly sampled from 1 minute long generated video, i.e. at 10, 20, 30, 40, 50, and 60 seconds. Our models can autoregressively generate 60-second, 1440-frame videos without quality degradation.
Full
Ablation 1
Ablation 2
Figure 6: Qualitative comparison for ablation study. Full represents for our full solution based on PA-M, Ablation 1 is with chunked frames but without overlapped conditioning. Ablation 2 is without both techniques. The frames are evenly sampled from 16-second generated videos.
Qualitative Results We also show strength of our method with qualitative comparison results in Fig. 5. Both of our models demonstrate strong performance in terms of frame fidelity and motion realism (e.g. camera motion, wave motion, and running gestures) and outperforms other baselines. For more qualitative results, please refer to our project webpage.
User study We conduct a human evaluation with 12 users to compare the generated videos from each method. As shown in Fig. 4, our PA-M is favored in each duel by a large margin.
4.3 Ablation Study
We conduct ablation studies on the PA-M model to evaluate the impact of chunked frames (Sec. 3.3), and overlapped conditioning (Sec. 3.4). Qualitative comparison is shown in Fig. 6 and in the project webpage. In Ablation 1, we observe that the absence of clean frames in the input sequence prevents noisy frames from attending to previous clean frames, resulting in poor performance over a long duration. This also causes frame-to-frame discontinuity, which is more noticeable in the project webpage. In Ablation 2, not decoding the video chunk-by-chunk leads to severe cumulative errors, causing the video to diverge after only a few seconds.
See Appendix H for additional ablation study on variable length and the number of sampling steps $S$.
5 Conclusion
In this work, we target long video generation, a fundamental challenge of current video diffusion models. We show that they can be naturally adapted to become progressive autoregressive video diffusion models without changing the architectures. With our progressive noise levels and the autoregressive video denoising process (Sec. 3.1), we achieve state-of-the-art results on 60-second long video generation. Since our method does not require model architecture changes, it can be seamlessly combined with orthogonal works, paving the way for generating longer videos at higher quality, long-term dependency, and controllability.
6 Acknowledgments
This research was supported in part by NSF award IIS2107224 and ONR award N000142312124.
Supplementary Material
Appendix A Summary
In this Appendix, we cover parallel works in Appendix B, related works in Appendix C, limitations and discussions in Appendix D, training details in Appendix E, evaluation details in Appendix F, additional qualitative results in Appendix G, and additional ablation study in Appendix H.
Appendix B Parallel Works
The core idea of our PA-VDM is to 1. assign progressively increasing noise levels to the $F$ frames in the attention window and 2. autoregressively apply the video diffusion model on progressively noised frames to generate long videos. The first part is inspired by Diffusion Forcing [4], which proposes to assign independent per-frame noise levels to some frames rather than a single noise level. We began developing our work right after July 1st, 2024, when Diffusion Forcing [4] was released on arXiv. The first version of our preprint was submitted to arXiv on October 10th, 2024. During this period, our work was developed independently, without the knowledge of two papers, Rolling Diffusion [38] and FIFO-Diffusion [19]. While Rolling Diffusion, FIFO-Diffusion, and PA-VDM arrive at a similar high-level idea in parallel, the three methods have different focuses, naming and framing of the idea, implementation details, experimental setups, and final result quality.
Compared to [38, 6], PA-VDM:
  1. 1.
    shows that it is possible to adapt a pre-trained video diffusion model to the progressive noise level schedule through finetuning, thus avoiding the otherwise immensely expensive computation cost of pre-training video diffusion models. [38] is trained from scratch and [19] is training-free.
  2. 2.
    achieves state-of-the-art 60-second long video generation at a quality comparable to frontier video diffusion models, demonstrating much longer video length and better quality than [38, 19].
We provide comparisons between our models (PA-M and PA-O) and [19] on our 60-second long video generation benchmark (Sec. F.2) in Secs. 4.2 and 1. Our method achieves substantially better qualitative and quantitative results than [19]. [19] also requires doubled inference cost, while our method only requires additional inference cost that is a fraction of the original cost (10% for PA-M and 16.66% for PA-O). We do not compare PA-VDM with [38] as there is no released code and it does not support text-conditioned open-domain generation.
Appendix C Related Works
The field of long video generation has faced significant challenges due to the computational complexity and resource constraints associated with training models on longer videos. As a result, most existing text-to-video diffusion models [10, 14, 15, 1] have been limited to generating fixed-size video clips, which leads to noticeable degradation in quality when attempting to generate longer videos. Recent works are proposed to address these challenges through innovative approaches that either extend existing models or introduce novel architectures and fusion methods.
Freenoise [36] utilizes sliding window temporal attention to ensure smooth transitions between video clips but falls short in maintaining global consistency across long video sequences. Gen-L-video [49], on the other hand, decomposes long videos into multiple short segments, decodes them in parallel using short video generation models, and later applies an optimization step to align the overlapping regions for continuity. FreeLong [28] introduces a sophisticated approach which balances the frequency distribution of long video features in different frequency during the denoising process. Vid-GPT [7] introduces GPT-style autoregressive causal generation for long videos.
More recently, Short-to-Long (S2L) approaches are proposed, where correlated short videos are firstly generated and then smoothly transit in-between to form coherent long videos. StreamingT2V [12] adopts this strategy by introducing the conditional attention and appearance preservation modules to capture content information from previous frames, ensuring consistency with the starting frames. It further enhances the visual coherence by blending shared noisy frames in overlapping regions, similar to the approach used by SEINE [5]. NUWA-XL [56] leverages a hierarchical diffusion model to generate long videos using a coarse-to-fine approach, progressing from sparse key frames to denser intermediate frames. However, it has only been evaluated on a cartoon video dataset rather than natural videos. VideoTetris [46] introduces decomposing prompts temporally and leveraging a spatio-temporal composing module for compositional video generation.
Another line of research focuses on controllable video generation [61, 45, 16, 60] and has proposed solutions for long video generation using overlapped window frames. These approaches condition diffusion models using both frames from previous windows and signals from the current window. While these methods demonstrate promising results in maintaining consistent appearances and motions, they are limited to their specific application domains which relies heavily on strong conditional inputs.
Appendix D Limitations and discussions
A limitation of our method is the demand of a well-trained base video diffusion model. Similar to the replacement methods [15, 58] and other approaches like StreamingT2V [12], our method autoregressively applies a video diffusion model to generate long videos. Such autoregressive video generation poses huge challenge on the base video diffusion model. Some slight errors remaining in the “clean” frames ${\mathbf{x}}^{0}$ may not be noticeable in a single video clip; however, in the autoregressive scenario, these error can be carried onto later frames, resulting in quality degradation. Further more, as the video diffusion model is only trained on denoising latent frames of real video data, it may poorly handle such distribution shift towards the generated erroneous frames [54, 6], resulting in more severe quality drop. This means that even after finetuning on our progressive noise levels, our method could still generate videos with some degree of quality degradation close to the ending, if the base video diffusion model is not well trained. Among the qualitative videos generated by our PA-M, in some cases, the video quality slightly degrades in the last 10 seconds.
Another limitation of our method is the subtle temporal flickering happening about every second in our PA-M results. It is caused by a flaw in the backbone video diffusion model M’s 3D VAE, as evident by the presence of such flickering in both PA-M and RW-M results while no such flickering is present in the PA-O results.
Appendix E Training details
M is pre-trained on captioned image and video datasets, containing 1 million videos and 2.3 billion images. These data are licensed and have been filtered to remove low-quality content. We train PA-M on video clips of $16,32,...,176$ raw frames that correspond to $F=5,10,...,55$ latent frames. The $F=55$ attention window length is derived by setting $F=S+5$, where $S=50$ is the number of sampling steps in M ($S=30$ in O) and $5$ is the length of an additional chunk of latent frames, as described in Secs. 3.3 and 3.4. The shorter latent frame lengths $F=5,10,...,50$ are used for the variable length training, as discussed in Sec. 3.2. RW-M is trained on videos of $64$ frames that corresponds to $F=20$ frames.
E.1 Modification to the base model
To implement progressive autoregressive video diffusion models on top of their pre-trained foundation video diffusion models, we do not need to modify the base model architectures. Instead, we only need to modify the model’s forward, training, and inference procedures. In the training and inference procedures, we replace the single noise level ${\textnormal{t}}\in[0,T)$ from regular diffusion model training [13, 15] with our per-frame noise level ${\mathbf{t}}_{0:F-1}$ and $\bm{\tau}^{\prime}_{0:S-1}$ (Secs. 3.5 and 3.1). To accommodate this change, we only need to make a single modification to the the noise level embedding computation in the model’s forward procedure. While the regular timestep only has the batch size dimension $B$, our progressive timesteps has two dimensions $B,F$. We first flatten them into the batch dimension of size $B\times F$, pass it to the timesteps embedding module, unflatten the two dimensions, and finally broadcast the timestep embedding to the same shape of the frames so they can be combined through either addition, concatenation, modulation, or cross-attention [33, 48, 32].
Appendix F Evaluation details
F.1 Baselines
As discussed in Sec. 4, using our base models, we implement two baseline autoregressive video generation methods on three models, which are denoted as RW-M, RN-O-base, and RN-O. We also compare to Stable Video Diffusion (SVD) [1] and StreamingT2V [12] model families. Specifically, we consider the SVD-XT model from SVD, a image-to-video model that generates a short video clip of 25 frames at 576x1024 resolution given an conditioning image. We apply it autoregressively, using the last image of the previous clip as the condition for generating a new clip. This is equivalent to the replacement-without-noise method except that it only conditions on a single frame rather than a chunk of 17 frames as RN-O. We also consider the StreamingSVD model from StreamingT2V, a image-to-long-video generation model that uses SVD as the base model [12]; its autoregressive video generation is enabled by training additional modules that connect to the base model via cross-attention. Similar to our progressive autoregressive video diffusion models, StreamingSVD can autoregressively generate long videos at 720x1280 resolution with arbitrary lengths, which we set to 1440 frames. We also compare to a concurrent work FIFO-Diffusion [19] implemented on Open-Sora-Plan v1.0.0 [23], denoted as FIFO-OSP. It generates at 256x256 resolution with a context window of 65 latent frames. See Appendix B for a discussion on [19] and other concurrent works. See Appendix F for details on our testing set, quantitative metrics, and traditional video quality evaluation.
FIFO-OSP
FIFO-Diffusion [19] is a parallel work that adopts a similar high-level idea as our method on pre-trained video diffusion models without any fine-tuning (see more discussion in Appendix B). It provides training-free implementations on VideoCrafter2 and Open-Sora-Plan v1.1.0 [23]. We choose its Open-Sora-Plan implementation since our method is also implemented on DiT-base [32] models, M and Open-Sora (O[58]. Open-Sora-Plan v1.1.0 generate videos at 512x512 resolution. Since there is no distributed inference support in the released code of FIFO-Diffusion, we adopt Open-Sora-Plan v1.0.0 in our reproduced FIFO-Diffusion results in order to saving computation costs by inferencing at the 256x256 resolution instead of the original 512x512 resolution.
F.2 Testing set
Text prompts and real videos
Our testing set consists of 40 text prompts and the corresponding real videos, sampled from Sora [58] demo videos, MiraData [18], UCF-101 [44], and LOVEU [52, 53]. For each text prompt, we generate two videos with 1440 frames, 60 seconds long at 24 FPS, resulting in a total of 80 videos. We use these 80 videos from each model for both quantitative and qualitative results, unless specified otherwise. Due to computation resource limitations of sampling 1-minute long videos, we only obtained partial results from M-PA, StreamingSVD and FIFO-OSP, including 48, 40, 40 videos from 24, 40, 40 text prompts respectively. This testing set measures the zero-shot long video generation ability of the models, since none of them are specifically trained on any of the above datasets.
Real video initialization
Since our focus is on long video generation, we focus on the video extension capability of the models rather than the text-to-short-video generation capability. Thus, we use the initial frames of the videos as the condition for all models, similar to the setting in [12]. M, O [58], StreamingSVD [12], SVD-XT [1], and FIFO-OSP [19, 23] use 16, 17, 1, 1, and 65 frames from the real video as the initial condition. Note that our PA-M and PA-O only require one chunk of frames (16 and 17 for M and O respectively), which is substantially less than the full context window of 65 frames required by FIFO-Diffusion [19]. This advantage is obtained from our variable-length autoregressive generation design as described in Sec. 3.2.
Appendix G Additional Qualitative Results
We provide additional quailtative results in Fig. 7.
@@ Appendix G Additional Qualitative Results
Figure 7: Qualitative comparison of PA-M (ours), RW-M, PA-O-base (ours), RN-O-base, StreamingSVD from StreamingT2V [12], SVD-XT from Stable Video Diffusion [1], and FIFO-Diffusion [19]. Frames are evenly sampled from 1 minute long generated video, i.e. at 10, 20, 30, 40, 50, and 60 seconds. Our models can autoregressively generate 60-second, 1440-frame videos without quality degradation.
Appendix H Additional Ablation Study
In our project webpage, we show an ablation study on our Variable Length design (Sec. 3.2). We compare Variable Length inference results of PA-M models trained with and without Variable Length. Without Variable Length training, the second video shows temporal jittering and abrupt scene change at the 1st and 59th seconds. This is because the model is not trained to generate the first/last chunk of latent frames to be consistent with the prior chunks. With Variable Length training, the first video avoids the jittering and abrupt scene change at the 1st and 59th seconds, and the video is temporally smooth. Furthermore, Variable Length inference enables the model to generate precisely 1440 frames, whereas without this technique the model would need to discard the noisy chunks remaining in the context window, which correspond to the 1441-1584th frames, when it reaches the 1440th frame. Being able to stop the autoregressive video denoising at a precise ending frame allows our model to generate a proper ending to the video, e.g. the woman exits the camera view in the first video, which is not possible without the Variable Length technique.
[…]
50 358.20
100 339.59
150 399.91
Table 2: Ablation on the number of sampling steps $S$ of the PA-M model.
Additionally, we ablate the number of sampling steps $S$ of the PA-M. Note that our progressive video denoising can work with arbitrary $S$; when the chunked frames technique is used, $S$ only needs to be divisible by $C$. We compute FVD scores in the same way as described in Sec. 4.2. As shown in Tab. 2, further increasing $S$ from 50 to 100 provides marginal benefits despite doubling the inference compute cost, while increasing $S$ to 150 leads to slightly worse results.